OpenStudy (anonymous):

Determine the midpoint between the points (4, 1) and (12, -3).

OpenStudy (lncognlto):

The formula for the midpoint is\[(\frac{ x1 + x2 }{ 2 }, \frac{ y1 + y2 }{ 2 })\] So, inserting the given values, we get\[(\frac{ 4 + 12 }{ 2 }, \frac{ 1 + (-3) }{ 2 })\]

OpenStudy (anonymous):

8, -1

OpenStudy (lncognlto):

Yes, that is correct! ^-^

OpenStudy (anonymous):

thanks sooo much

OpenStudy (lncognlto):

You're welcome :)